OpenStudy (alexwee123):

In a school girl's heights are normally distributed w/ a mean of 5'4" and a standard deviation of 2". a)What is the probability of a randomly chosen girl being taller than 5'5"? b) If the principal chooses 5 girls at random, what is the probability that at least 4 of them are tallker than 5'5"?

OpenStudy (alexwee123):

i got 52% for A)

OpenStudy (anonymous):

damn this is a higher level of probability, and i totally forgot how to do this kind of questions. Gimme some time i go refresh my memory. Btw are you singaporean?

OpenStudy (alexwee123):

lol no i'm korean american :)

OpenStudy (anonymous):

Okay, first you must understand that the heights are normally distributed. so calculate the z-value to get the distribution of those values in a normal distribution curve. (in simpler terms, to calculate z-value means to adjust the values given to fit in the standard normal distribution curve) So the z-value for being SHORTER than or EQUAL to 5'5" = (5'5" - 55'4") / 2" = 0.5 The associated probability for z value of 0.5 is 0.70 ( http://www.math.unb.ca/~knight/utility/NormTble.htm) Since you want the probability of the girl TALLER than 5'5", 1- 0.7 = 0.3 b) If the principal chooses 5 girls at random, what is the probability that at least 4 of them are tallker than 5'5"? P ( at least 4 girls taller than 5"5") = P (4 girls taller than 5'5") + P( 5 girls taller than 5'5") = 0.3*0.3*0.3*0.3*0.7 + 0.3*0.3*0.3*0.3*0.3= 0.0081

OpenStudy (anonymous):

z score = ( 5.5 - 5.4 ) / 2 = .05 => P ( .05 ) = .4801

OpenStudy (alexwee123):

shouldn't you convert 5'5" and 5'4" into inches since the standard deviation is in inches?

OpenStudy (alexwee123):

@tiaph why is it P (4 girls taller than 5'5") + P( 5 girls taller than 5'5") ?

OpenStudy (anonymous):

I didnt convert because, 5ft5 inches - 5ft4inches = 1 inch and 1 inch /2 inch = 0.5

OpenStudy (anonymous):

Cause the question ask for at least 4 girls taller than 5'5". meaning 2 situations which you have to take into account. Situation 1: 4 girls taller than 5'5" and 1 girl shorter or equal to 5'5" Situation 2: all 5 girls taller than 5'5"

OpenStudy (alexwee123):

ohhhhhh i remember now thanks:D!

OpenStudy (alexwee123):

and if the question is asking if "at most" 4 girls then it would be P(1 girl taller) +P(2girls taller)....+P(4girls taller) right?

OpenStudy (anonymous):

yes you got it ;D

OpenStudy (alexwee123):

w8 shouldn't it be 0.7x0.7x0.7x0.3 +0.7^5 because it's asking for the probability that 4 are taller not shorter?

OpenStudy (anonymous):

0.7 is the probability for the girls SHORTER or EQUAL to 5'5". If you want taller 0.3 is the answer you are looking for. Don't get confused.

OpenStudy (alexwee123):

but if you look at it